# MultiAPI Spoof: A Multi-API Dataset and Local-Attention Network for Speech Anti-spoofing Detection thanks: Corresponding Author: Ming Li

### v1

@@ 1 Introduction
Recently, text-to-speech (TTS) [tts1, tts2, tts3, tts4], voice conversion (VC) [vc1, vc2, vc3, vc4], and generative modeling techniques have evolved rapidly. In particular, end-to-end dialogue systems [Qwen3-omni, Baichuan-omni], speech continuation models [Baichuan-audio], and style- or emotion-specific speech generation models [SoulXPodcast, Step-Audio-EditX] have advanced significantly. As a result, synthetic speech has become increasingly realistic and pervasive in everyday applications. Modern audio generation systems, especially those based on diffusion and large-scale generative models [diffu, Qwen3-omni, Kimi-audio], can now produce speech that closely mimics human prosody, timbre, and emotion. However, they have introduced serious security risks and can be easily misused for impersonation or misinformation.
@@
To address these limitations, we introduce MultiAPI Spoof, a new multi-API speech anti-spoofing dataset designed for both anti-spoofing detection and API-level source tracing. Unlike prior datasets that focus on a few synthesis systems, MultiAPI Spoof comprises audio generated from 30 distinct APIs, including commercial TTS services, open-source speech models, and TTS websites. The dataset covers approximately 230 hours of spoofed speech. Based on the MultiAPI Spoof dataset, (1) we show that there is a gap between previous research benchmarks and real-world spoofing scenarios; adding our API dataset in the training can also enhance the performance on current benchmarks. (2) Furthermore, we propose a new detection method, namely Nes2Net-LA, built upon Nes2Net [nes2net]. By integrating local attention modules between Nested blocks, Nes2Net-LA enhances local context modeling and fine-grained spoofing feature extraction, thereby improving robustness and discriminative capability. The Nes2Net-LA achieves state-of-the-art (SOTA) performance across multiple anti-spoofing benchmarks. (3) Finally, we introduce the API tracing task, which aims to identify the generation API of spoofed audio and establishes a benchmark for fine-grained source attribution.
@@ 3.1 Prior Knowledge: Nested Res2Net (Nes2Net-X)
wo
MultiAPI
Spoof
XLSR+AASIST [xlsr+aasist] 2.02 / 0.026 / 0.029 - - 7.30 / 0.098 / 0.106 12.96 / 0.132 / 0.190
@@
with
MultiAPI
Spoof
XLSR+AASIST [xlsr+aasist] 2.09 / 0.028 / 0.030 0.48 / 0.007 / 0.0070 0.83 / 0.010 / 0.012 0.70 / 0.009 / 0.010   6.26 / 0.079 / 0.092
@@ 5.2.2 Effectiveness of Local Attention (Nes2Net-LA)
As shown in Table 1 and Table 2, Nes2Net-LA trained on the data collection pool outperforms recent state-of-the-art models across all evaluation benchmarks, even without any data augmentation or pruning. The most substantial improvements are observed on the unseen split of the MultiAPI Spoof test set. These results demonstrate that the proposed local attention mechanism effectively enhances cross-block interactions and produces more discriminative and robust anti-spoofing representations.

### v2

@@ 1 Introduction
Recently, Text-To-Speech (TTS) [tts1, tts2, tts3, tts4], Voice Conversion (VC) [vc1, vc2, vc3, vc4], and generative modeling techniques [qwen2audio, gpa, gpt4, Gemini] have evolved rapidly. In particular, end-to-end dialogue systems [fun, Qwen3-omni, wavbench, Baichuan-omni], speech continuation models [Baichuan-audio], and style- or emotion-specific speech generation models [SoulXPodcast, flexivoice, Step-Audio-EditX] have advanced significantly. As a result, synthetic speech has become increasingly realistic and pervasive in everyday applications. Modern audio generation systems, especially those based on diffusion and large-scale generative models [diffu, Qwen3-omni, Kimi-audio], can now produce speech that closely mimics human prosody, timbre, and emotion. However, they have introduced serious security risks and can be easily misused for impersonation or misinformation.
@@
To address these limitations, we introduce MultiAPI Spoof, a new multi-API speech anti-spoofing dataset designed for both anti-spoofing detection and API-level source tracing. Unlike prior datasets that focus on a few synthesis systems, MultiAPI Spoof comprises audio generated from 30 distinct APIs, including commercial TTS services, open-source speech models, and TTS websites. The dataset covers approximately 230 hours of spoofed speech. Based on the MultiAPI Spoof dataset, our contributions are as follows:
  1. 1.
    We show that there is a gap between previous research benchmarks and real-world spoofing scenarios; adding our API dataset in the training can also enhance the performance on current benchmarks.
  2. 2.
    Furthermore, we propose a new detection method, namely Nes2Net-LA, built upon Nes2Net [nes2net]. By integrating local attention modules between Nested blocks, Nes2Net-LA enhances local context modeling and fine-grained spoofing feature extraction, thereby improving robustness and discriminative capability. The Nes2Net-LA achieves state-of-the-art (SOTA) performance across multiple anti-spoofing benchmarks.
  3. 3.
    Finally, we introduce the API tracing task, which aims to identify the generation API of spoofed audio and establishes a benchmark for fine-grained source attribution.
@@ 3.1 Prior Knowledge: Nested Res2Net (Nes2Net-X)
wo MultiAPI Spoof XLSR+AASIST [xlsr+aasist] 2.02 / 0.026 / 0.029 - - 7.30 / 0.098 / 0.106 12.96 / 0.132 / 0.190
@@
with MultiAPI Spoof XLSR+AASIST [xlsr+aasist] 2.09 / 0.028 / 0.030 0.48 / 0.007 / 0.0070 0.83 / 0.010 / 0.012 0.70 / 0.009 / 0.010   6.26 / 0.079 / 0.092
@@ 5.2.2 Effectiveness of Local Attention (Nes2Net-LA)
As shown in Table 1 and Table 2, Nes2Net-LA trained on the data collection pool outperforms recent state-of-the-art models across all evaluation benchmarks, even without any data augmentation or pruning. The most substantial improvements are observed on the unseen split of the MultiAPI Spoof test set, demonstrating that the proposed local attention mechanism effectively enhances cross-block interactions and produces more discriminative and robust anti-spoofing representations.